Lauran Smith is planning for her and her husband Luke’s retirement. Both Luke and Laura expect to retire in 35 years (when they turn 65). The life expectancy of men is 75 years and the life expectancy of women is 85 years (i.e., assume that they die the day before their 75 th or 85 th birthday). During retirement (while they are living), the couple wants to withdraw $10,000 at the beginning of each year from their savings account- $5,000 for each of them. Assume that the interest rate during their retirement is 10 percent compounded annually; the interest rate after Luke dies is 12% compounded semi-annually; and, the interest rate prior to retirement is 9 percent compounded annually . How much will they have to deposit in their joint savings account each month (beginning one month from now and ending on their retirement date)?

(b) How would your answer change if Luke and Laura Smith expect to inherit $40,000, 5 years from now? Assume that, at that time, they will spend $10,000 for a two-week luxury all-inclusive cruise , spend $25,000 to buy top-of-the-line stainless steel appliances for their kitchen, and save the rest for their retirement. Assume equal monthly payments throughout the 35 years to retirement- i.e., the same monthly payments beginning one month from now and ending on their retirement date.
Answer #1
Inputs:
Years to retirement 35
Months to retirement =35*12= 420
Length of retirement (in years) for Luke 10 (75-65)
Length of retirement (in years) for Lauran Smith 20 (85-65)
Annual interest rate for first 10 years of retirement 10% 1.007207323
Semi annual interest for last 10 years of retirement 6% (12/2)
Annual interest for last 10 years of retirement=(1.06^2)-1 12.36%
Annual interest rate prior to retirement 9%
Monthly interest rate prior to retirement =r
(1+r)^12=1+0.09=1.09, 1+r=1.09^(1/12)=1.007207323
Monthly interest rate prior to retirement =r 0.7207%
Withdrawal for first 10 years after retirement(beginning of year) $10,000
Withdrawal for last 10 years after retirement(beginning of year) $5,000
Type for beginning of period withdrawal                        1
Ouputs:
A Present Value of first 10 years withdrawal at the time of retirement $67,590 (Using PV Function of excel with Rate=10%,Nper=10,Pmt=-10000,Type=1)
Present Value of Last 10 years withdrawal at the time death of Luke $31,281 (Using PV Function of excel with Rate=12.36%,Nper=10,Pmt=-5000, Type=1)
B Present Value of Last 10 years withdrawal at the time of retirement $12,060 (31281/(1.1^10)
C=A+B            Total amount needed at retirement $79,650
D Required Savings Per MONTH for 35 Years $29.57 (Using PMT Function of excel with Rate=0.7207%,Nper=420,Fv=-79650)
b Additional inputs
Inheritance after 5 years $40,000
Amount spent from inheritance=$10000+$25000 $35,000
Net Saving after 5 years $5,000 (40000-35000)
Number of years the savings earns 9% annual interest 30 (35-5)
E Future Value of savings of $5000 at the time of retirement $66,338 (Using FV Function of excel with Rate=9%,Nper=30,Pv=-5000))
F=C-E SHORTFALL   $13,312 (79650-66338)
Required Savings Per MONTH for 35 Years $4.94 (Using PMT Function of excel with Rate=0.7207%,Nper=420,Fv=-13312)

LIPDUdlu PUPIL Allymeril Number Styles Cells LUILITY H21 X Fac =PV(H9, H7,-H16,,H18) H i I K L M N O 6 Months to retirement =LIPDUdlu PUPIL Allgrimeril Number Styles Cells LUILITY H22 X fac =PV(H11,10,-H17,,H18) H I J K L M N O 6 Months to retirementH25 X fo PMT(H15, H6,,-H24) F L M N O Semi annual interest for last 10 years of retirement 6%|(12/2) Annual interest for lastH35 X V for PMT(H15,H6, -H34) H I J K L M N O 21 A $67,590 (Using PV Function of excel with Rate=10%,Nper=10, Pmt=-10000, Typ